# NormTab: Improving Symbolic Reasoning in LLMs Through Tabular Data Normalization

### v1

@@ 4.3 Implementation
We utilized gpt-3.5-turbo-0125 as the Language Model which supports 16k context window. We were inspired by the prompting style from (Liu et al., 2023; Nahid and Rafiei, 2024) in our implementation of NormTab. To compare performance, we employ few-shot in-context learning. This involves supplying the LLM with the table title, table header, question, and three example rows of the table, along with the question, to generate an SQL query. The SQL query is then executed on the table to obtain the answer. Further details can be found in Appendix C  [^1] The source code for the implementation discussed in this paper will be made publicly available upon acceptance of the paper..
@@ 6 Conclusion
Despite the advancements brought by NormTab, there are several limitations. First, while our framework significantly enhances the symbolic reasoning capabilities of LLMs on tabular data, there remains room for improvement in the normalization process, particularly with more complex table structures. Additionally, for larger tables, LLMs may sometimes produce hallucinated results, leading to inaccuracies in the normalized output, indicating a need for better handling of extensive datasets. Furthermore, when dealing with tables that contain extremely noisy data, LLMs struggle to effectively clean and normalize the information. The presence of excessive noise and inconsistencies can hinder the normalization process and negatively impact overall performance. Addressing these limitations is crucial for further enhancing the robustness and reliability of NormTab. As we measure the accuracy using the results obtained from LLM based Text-to-SQL model, it is important to note that some questions in the dataset may not directly map to SQL queries which may affect the performance.
@@ Appendix B Comparison with Other Models
Appendix C Implementation
We configured the in-context learning hyperparameters for gpt-3.5-turbo-0125 according to the specifications outlined in Table 6.

### v2

@@ 4.3 Implementation
We utilized gpt-3.5-turbo-0125 as the Language Model which supports 16k context window. We were inspired by the prompting style from (Liu et al., 2023; Nahid and Rafiei, 2024) in our implementation of NormTab. To compare performance, we employ few-shot in-context learning. This involves supplying the LLM with the table title, table header, question, and three example rows of the table, along with the question, to generate an SQL query. The SQL query is then executed on the table to obtain the answer. Further details can be found in Appendix C, and all our code and prompts are available at https://github.com/mahadi-nahid/NormTab.
@@ 5.1 Results on Downstream Tasks
In this study, we also evaluated the effectiveness of our method using Gemini-1.5-flash and GPT-4-turbo. The results showed a improvement on the WikiTableQuestions dataset, demonstrating that our model’s performance is not heavily dependent on specific language models (see Table 3).
Model Acc (WTQ)
NormTabBasic (gemini-1.5-flash) 61.36
NormTabTargeted (gemini-1.5-flash) 61.24
NormTabBasic (gpt-4-turbo) 61.57
NormTabTargeted (gpt-4-turbo) 62.28
Table 3: Performance of NormTab on WikiTQ dataset using Gemini-1.5-flash and GPT-4-turbo.
@@ 5.2 NormTab Evaluation
Our normalization method, NormTab, can be beneficial for a variety of table reasoning tasks, especially those employing symbolic methods. For the same reason, we have integrated NormTab with a recent table reasoning method TabSQLify (Nahid and Rafiei, 2024), which utilizes symbolic techniques. Our evaluation reveals that integrating NormTab with TabSQLify leads to a notable 4% improvement in performance. This demonstrates NormTab’s potential to enhance other symbolic frameworks by serving as an effective preprocessing step. The results is summarized in Table 5.
Model Acc (WTQ)
TabSQLify 64.7
NormTab+TabSQLify 68.63
Table 5: Performance of TabSQLify integrated with NormTab on the WikiTableQuestions dataset (gpt-3.5-turbo).
In our work, each table is normalized once, as a preprocessing step, irrespective of the number and the type of questions asked. Our approach does not depend on the specific question or task. While using Self-Consistency (Wang et al., 2023; Liu et al., 2023) might seem beneficial, it requires generating multiple responses per question, which can increase costs and reduce efficiency. For instance, if we need answers to 10 questions and apply Self-Consistency with 6 paths, we end up with 60 samples (10 questions × 6 paths). Additionally, we need to send the entire table with each question, resulting in a higher number of tokens. In contrast, our method requires just a few prompts to normalize the table initially. After normalization, we do not need to pass the entire table to generate the SQL query for each question; only table metadata such as the table title, column names, and a few example rows, is needed. This one-time preprossessing step significantly reduces the overall number of tokens passed to the LLM, potentially lowering costs compared to using Self-Consistency.
@@ 5.3 Analysis
The basic NormTab approach involves only one LLM call, but it requires passing the entire table to the language model as a prompt. This means the LLM must process a larger number of tokens, which can impact the overall normalization performance. Research indicates that more tokens can increase the likelihood of hallucination and can be more costly (Chen, 2023; Ye et al., 2023; Ji et al., 2023). In contrast, the targeted NormTab approach first filters out the parts of the table that are already well-formatted, thereby reducing the number of tokens sent to the LLM by focusing only on the columns that need normalization. For example, consider a table with 15 rows and 8 columns, resulting in a total of 15 * 8 = 120 table cells. In the basic NormTab approach, all 120 cells are sent to the language model along with the instructions. However, if we identify that only 5 out of 8 columns require normalization, we only need to send 15 * 5 = 75 table cells. This reduction translates to 45 fewer table cells, which is a 37.5% reduction in table size. Table 8 illustrates that we can achieve a 72% reduction in table size for both datasets by employing the targeted NormTab approach, which is quite substantial.
Dataset
NormTab
NormTab
Reduction
WikiTQ 152.26 41.82 72.53%
TabFact 106.19 29.11 72.58%
Table 8: Table cell reduction in NormTab-Targeted compared to NormTab-Basic
Although the targeted NormTab approach requires additional LLM calls for tasks such as column selection and transposition detection, the total number of tokens processed is still lower than in the basic NormTab approach. While the basic approach may be suitable for smaller tables, the reduction in table size is crucial for normalizing larger tables effectively. The targeted strategy involves multiple queries, but it refines the normalization process by concentrating on specific subtasks, which may help reduce hallucination and errors. While the performance improvement appears marginal, the significant token size reduction makes the targeted NormTab approach highly beneficial for larger tables.
@@ 6 Conclusion
Despite the advancements brought by NormTab, there are several limitations. First, while our framework significantly enhances the symbolic reasoning capabilities of LLMs on tabular data, there remains room for improvement in the normalization process, particularly with more complex table structures. Additionally, for larger tables, LLMs may sometimes produce hallucinated results, leading to inaccuracies in the normalized output, indicating a need for better handling of extensive datasets. Moreover, when working with tables containing highly noisy data, LLMs often struggle to clean and normalize the information effectively, and may generate output in an incorrect format, making it challenging to parse. The presence of excessive noise and inconsistencies can hinder the normalization process and negatively impact overall performance. Addressing these limitations is crucial for further enhancing the robustness and reliability of NormTab. As we measure the accuracy using the results obtained from LLM based Text-to-SQL model, it is important to note that some questions in the dataset may not directly map to SQL queries which may affect the performance.
Acknowledgments
We extend our sincere gratitude to all reviewers for their invaluable feedback, insightful suggestions, and positive remarks about our work. This research has been supported by the Natural Sciences and Engineering Research Council of Canada. Also, Md Mahadi Hasan Nahid was supported by the Alberta Innovates Graduate Student Scholarship.
@@ Appendix B Comparison with Other Models
Appendix C Implementation Settings
The dataset we used contains only the gold answers and lacks the original SQL queries needed to extract these answers. Additionally, the dataset does not include normalized or cleaned versions of the tables. Data contamination can indeed impact methods where the question and table are provided to the LLM, which might then rely on knowing the gold answer directly. In our approach, however, we focus on converting tables into a normalized format. Since the dataset does not provide such cleaned versions, data contamination does not affect our method. Moreover, our approach generates SQL queries based on the question and table metadata, rather than relying on pre-existing gold SQL queries. We obtain the answers by executing these generated SQL queries. Therefore, data contamination does not impact our method.
Our implementation of NormTab was inspired by the prompting techniques used in (Liu et al., 2023; Nahid and Rafiei, 2024). We configured the in-context learning hyperparameters for gpt-3.5-turbo-0125 according to the specifications outlined in Table 9.